90 kN/m 165 KN 120 KN.m ܠܠܠܠ 2m 3m 1m 1m А в 2m С DE F

which among the choices almost give the reaction at the fix support? use any method

A. 8.312
B. 52.660
C. 255.653
D. 85.427
Answer #1

1. EMC F 2 =G 2 Ey 120 8 By + 2 + ME +16589 +[(4 83x90) *4] 8 By + 2 Ey = 2145 Eq.(1) + ME = LEY By + Ey + Fy 165 +( *3*90) •Integrate 2 EI dy - y <x) 2) ME <x> + Ey <x-27 !!! dx 2) + By <*-»> +120 64-6 2) G + 165 <x-gy 21 go <x-67 3 4) +90 <x-37 3x=0 At No downward deflection from Fq? ci) C2 Also, At x = 2 m ; y=0 Eq? co Becomes EI (o) Fy (2) - MF (2) 2 2) 3) 1 - 3333 Fje; o = 85.3333 Fy 32 ME +36EY + 240 8 + 781.25 +2343.75 j.e. 85.3333 Fy -32 MF +36 Ey = - 3357 Eq. (4) We have 4 Eq. (I to 4